6 22 + 3 6) Of all lines tangent to the graph of y = maximum slope. find the tangent lines of minimum slope and
Answer #1

The given function is f(x) = y= 6 22 +3 f(x) = 12x (x2+3)2 . g(x) = f(x) = - Let 12x (x2+3)2 When the tangent line has the介 The value of the derivative of the function f(x) is maximum at x = -1 and minimum at x = 1. 6 When X = -1, y = 1+3 3 When x